### v1

# Gabor systems with Hermite functions of order $n$ and oversampling greater than $n+1$ which are not frames

Abstract.
We show that a sufficient density condition for Gabor systems with Hermite functions over lattices is not sufficient in general. This follows from a result on how zeros of the Zak transform determine the frame property of integer over-sampled Gabor systems.
Key words and phrases:
Gabor frame, Hermite function, periodic configuration
1991 Mathematics Subject Classification:
42C15; 33C45
This research was funded in whole or in part by the Austrian Science Fund (FWF) [10.55776/P33217].
1. Introduction
Gabor systems with Hermite functions have some quite satisfying, but just as many mysterious properties. A fundamental result due to Lyubarskii [11] and Seip and Wallstén [12][13] is that a Gabor system with a Gaussian window is a frame for $L^{2}({\mathbb{R}})$ if and only if the sampling rate exceeds 1. Indexing the Gaussian as the 00-th Hermite function, we have the quite satisfying condition, due to Gröchenig and Lyubarskii [6], [7], that a Gabor system with the $n$-th Hermite function over a lattice $\Lambda\subset{\mathbb{R}}^{2}$ is a frame once the density of the lattice is larger than $n+1$, i.e., ${\rm vol}({\mathbb{R}}^{2}/\Lambda)<1/(n+1)$. A main difference to the results in [11], [12], [13] is that the results in [6] and [7] rely on the lattice structure.
Recently, Gabor frames with Hermite functions and oversampling rate $n+1$ were found by Faulhuber, Shafkulovska, and Zlotnikov [2]. On the other side, simple conditions for the frame property of Gabor systems with Hermite functions are to be ruled out eventually, due to the results of Lemvig [10] or Horst, Lemvig, and Videbaek [8].
In this work, we are concerned with the following question: can the sufficiency condition of Gröchenig and Lyubarskii be extended beyond the lattice setting? The answer is NO!
2. Notation
In most parts of this article we will follow the notation from the book of Gröchenig [5]. The Hilbert space of square-integrable functions on the real line is denoted by $L^{2}({\mathbb{R}})$. For technical reasons, it will sometime be necessary to restrict ourselves to a suitable dense subspace, in which case we will also write that a function is suitable. The statement is then valid for the Hilbert space by the density of the suitable function space, with the constraint that some results will only hold almost everywhere and not pointwise.
The modulation space $M^{1}({\mathbb{R}})$, also known as the Feichtinger algebra $S_{0}({\mathbb{R}})$ [3], is a suitable function space. In some situations, functions from the (slightly larger) Wiener space of continuous functions $W_{0}({\mathbb{R}})$ with Fourier transform also contained in $W_{0}({\mathbb{R}})$ suffice. In any case, our results will be for Hermite functions, which are contained in the Schwartz space $\mathcal{S}({\mathbb{R}})$ and the reader may consider their preferred function space as suitable.
The Fourier transform of a suitable function on ${\mathbb{R}}$ is given by
[…]
Note that $\mathcal{F}$ extends to a unitary operator on $L^{2}({\mathbb{R}})$, i.e.,
[…]
Moreover, we will use the unitary dilation operator $\mathcal{D}_{a}$, $a>0$, which acts on suitable functions by the rule
(2.3) $\mathcal{D}_{a}f(t)=\frac{1}{\sqrt{a}}f\left(\frac{t}{a}\right).$
This clarifies the basic essentials. In § 2.1 we clarify our normalization for Hermite functions, which is the same as in the book of Folland [4]. Time-frequency shifts and multi-window Gabor systems are introduced in § 2.2, following the book of Gröchenig [5]. Also, the notion of a periodic structure (the union of finitely many shifted copies of one lattice) is discussed there. We then introduce the (parameter free) Zak transform in § 2.3, which maps functions from ${\mathbb{R}}$ to quasi-periodic functions with period lattice ${\mathbb{Z}}^{2}$. This is again covered by the book of Gröchenig [5]. The reader familiar with the concepts and notation may directly jump to the results in § 3.
2.1. Hermite functions
We consider Hermite functions as defined in [4]. The $n$-th order Hermite function is the $n$-th Hermite polynomial $H_{n}$ (subject to our normalization) multiplied with a Gaussian function $\phi(t)=e^{-\pi t^{2}}$. More precisely, we have
(2.4) $h_{n}(t)=(-1)^{n}C_{n}e^{\pi t^{2}}\dfrac{d^{n}}{dt^{n}}\left(e^{-2\pi t^{2}}\right)=H_{n}(t)\phi(t),$
where $C_{n}$ is a normalizing constant, which is explicitly given by $C_{n}=2^{1/4}/\sqrt{n!(2\pi)^{n}2^{n}}$. In particular, for $n=0$ we obtain a normalized Gaussian function
(2.5) $h_{0}(t)=C_{0}\,\phi(t)=2^{1/4}e^{-\pi t^{2}}.$
Moreover, the Hermite functions are eigenfunctions of the Fourier transform with
(2.6) $\mathcal{F}h_{n}=(-i)^{n}h_{n},\quad n\in{\mathbb{N}}_{0}.$
2.2. Gabor systems
A Gabor system in $L^{2}({\mathbb{R}})$ is a function system of the type
(2.7) $\mathcal{G}(g,\Gamma)=\{\pi(\gamma)g\mid\gamma\in\Gamma\subset{\mathbb{R}}^{2}\},$
where $\pi(\gamma)$ is a time-frequency shift. This is the composition of a translation (time-shift) and a modulation (frequency-shift), i.e.,
(2.8) $\pi(z)=M_{\omega}T_{x},\ z=(x,\omega)\quad T_{x}f(t)=f(t-x),\quad M_{\omega}f(t)=e^{2\pi i\omega t}.$
Time-frequency shifts are not closed under composition due to the commutation relation
(2.9) $M_{\omega}T_{x}=e^{2\pi i\omega x}T_{x}M_{\omega}.$
However, by adding a phase factor $c\in\mathbb{C}$, $|c|=1$, they become a non-commutative group:
(2.10) $e^{2\pi i\tau_{1}}\,\pi(z_{1})\,e^{2\pi i\tau_{2}}\,\pi(z_{2})=e^{2\pi i(\tau_{1}+\tau_{2}-x_{1}\omega_{2})}\,\pi(z_{1}+z_{2}),\quad z_{1}=(x_{1},\omega_{1}),\,z_{2}=(x_{2},\omega_{2}).$
We refer to [5, Chap. 9.1] for more details. The important part for us here is the fact that the composition of two time-frequency shifts is (up to a phase factor) again a time-frequency shift by the sum of the co-ordinates.
A Gabor system is a frame for $L^{2}({\mathbb{R}})$ if and only if there exist constants $0<A\leq B<\infty$, called frame bounds, such that the following frame inequality is satisfied:
(2.11) $A\lVert f\rVert_{L^{2}}^{2}\leq\sum_{\gamma\in\Gamma}|\langle f,\pi(\gamma)g\rangle|^{2}\leq B\lVert f\rVert_{L^{2}}^{2},\quad\forall f\in L^{2}({\mathbb{R}}).$
The optimal frame bounds are the spectral bounds of the associated Gabor frame operator
(2.12) $S_{\mathcal{G}(g,\Gamma)}f=\sum_{\gamma\in\Gamma}\langle f,\pi(\gamma)g\rangle\,\pi(\gamma)g.$
We remark that we always assume $\Gamma$ to be relatively separated, i.e.,
(2.13) $\min_{\begin{subarray}{c}\gamma_{1},\gamma_{2}\in\Gamma\\ \gamma_{1}\neq\gamma_{2}\end{subarray}}(|\gamma_{1}-\gamma_{2}|)\geq\delta>0.$
In this case, for $g\in M^{1}({\mathbb{R}})$ the upper frame bound is always finite and we only need to care about the strict positivity of the lower frame bound.
A multi-window Gabor system is of the form
(2.14) $\mathfrak{G}_{M}(g_{m},\Gamma)=\bigcup_{m=1}^{M}\mathcal{G}(g_{m},\Gamma)$
and the associated multi-window frame operator is given by
(2.15) $S_{\mathfrak{G}_{M}(g_{m},\Gamma)}=\sum_{m=1}^{M}S_{\mathcal{G}(g_{m},\Gamma)}=\sum_{m=1}^{M}\sum_{\gamma\in\Gamma}\langle f,\pi(\gamma)g_{m}\rangle\,\pi(\gamma)g_{m}.$
The windows $g_{m}$ may be completely different, but we are interested in the case where they are simply shifted copies $\pi(z_{m})g$ of a single function $g$. A direct computation shows that the multi-window frame operator is also the associated frame operator of a Gabor system over a periodic structure $\Gamma_{M}(\Lambda,\{z_{m}\}_{m=1}^{M})$ with window $g$. The periodic structure is
(2.16) $\Gamma_{M}(\Lambda,\{z_{m}\}_{m=1}^{M})=\bigcup\limits_{m=1}^{M}(\Lambda+z_{m})$
To simplify notation, we will only write $\Gamma_{M}(\Lambda)$, keeping in mind that we still need to choose the set $\{z_{1},\ldots,z_{m}\}$. While $\Gamma_{M}(\Lambda)$ is still $\Lambda$-periodic it is not necessarily a lattice. The density, or (over-)sampling rate, of a periodic structure is given by
(2.17) $\rho(\Gamma_{M}(\Lambda))=\frac{M}{{\rm vol}({\mathbb{R}}^{2}/\Lambda)}.$
Assume $g$ is suitably nice (so the following series converges unconditionally), then
(2.18) $S_{\mathfrak{G}_{M}(g_{m},\Gamma_{M}(\Lambda))}f$ $=\sum_{m=1}^{M}\sum_{\lambda\in\Lambda}\langle f,\pi(\lambda)\pi(z_{m})g\rangle\,\pi(\lambda)\pi(z_{m})g$
(2.19) $=\sum_{m=1}^{M}\sum_{\lambda\in\Lambda}\langle f,\pi(\lambda+z_{m})g\rangle\,\pi(\lambda+z_{m})g=S_{\mathcal{G}(g,\Gamma_{M})}f.$
The phase factor, which appears due to (2.9), also appears as complex conjugate in the above calculation and, hence, cancels. The last detail we collect is the unitary equivalence of the Gabor systems $\mathcal{G}(g,\Gamma)$ and $\mathcal{G}(\mathcal{D}_{a}g,D_{a}\Gamma)$, where
(2.20) $D_{a}=\begin{pmatrix}a&0\\ 0&\frac{1}{a}\end{pmatrix},\quad a>0.$
First, for $z\in{\mathbb{R}}^{2}$, we note that $\mathcal{D}_{a}\,\pi(z)\,\mathcal{D}_{a}^{-1}=\pi(D_{a}z)$, which can be checked by a simple calculation with a suitable function. From this we get
(2.21) $S_{\mathcal{G}(\mathcal{D}_{a}g,D_{a}\Gamma)}f$ $=\sum_{\gamma\in D_{a}\Gamma}\langle f,\pi(\gamma)\mathcal{D}_{a}g\rangle\,\pi(\gamma)\mathcal{D}_{a}g=\sum_{\gamma\in\Gamma}\langle f,\pi(D_{a}\gamma)\mathcal{D}_{a}g\rangle\,\pi(D_{a}\gamma)\mathcal{D}_{a}g$
(2.22) $=\sum_{\gamma\in\Gamma}\langle f,\mathcal{D}_{a}\pi(\gamma)\mathcal{D}_{a}^{-1}\mathcal{D}_{a}g\rangle\,\mathcal{D}_{a}\pi(\gamma)\mathcal{D}_{a}^{-1}\mathcal{D}_{a}g=\mathcal{D}_{a}\sum_{\gamma\in\Gamma}\langle\mathcal{D}_{a}^{-1}f,\pi(\gamma)g\rangle\,\pi(\gamma)g$
(2.23) $=\mathcal{D}_{a}S_{\mathcal{G}(g,\Gamma)}\mathcal{D}_{a}^{-1}f.$
In the above computations we also used the unconditional convergence of the series (because $g$ is suitable, so we have a Bessel system, i.e., a finite upper frame bound) and the fact that $\mathcal{D}_{a}$ is unitary. In particular, the two Gabor systems have the same optimal frame bounds.
2.3. The Zak transform
The Zak transform has become a standard tool in time-frequency analysis to study the frame property of integer over-sampled Gabor systems. For a suitable function $f$, its Zak transform is given by
(2.24) $\mathcal{Z}f(x,\omega)=\sum_{k\in{\mathbb{Z}}}f(k-x)e^{2\pi ik\omega}=\sum_{k\in{\mathbb{Z}}}M_{\omega}T_{x}f(k).$
In shorter notation, setting $z=(x,\omega)$ we can also write this as
(2.25) $\mathcal{Z}f(z)=\sum_{k\in{\mathbb{Z}}}\pi(z)f(k).$
We note the following basic properties, which are easily verified.
(2.26) $\mathcal{Z}f(x+1,\omega)=e^{2\pi i\omega}\mathcal{Z}f(x,\omega)\quad\text{ and}\quad\mathcal{Z}f(x,\omega+1)=\mathcal{Z}f(x,\omega)$
Thus, $\mathcal{Z}f(x,\omega)$ is quasi-periodic with respect to the integer lattice ${\mathbb{Z}}^{2}$ and, thus, completely determined by the values in $[0,1)^{2}$. Next, we note that a time-frequency shift of the function basically results in a shift of the Zak transform. Due to (2.9), we obtain
(2.27) $\mathcal{Z}(M_{\eta}T_{\xi}f)(x,\omega)=e^{-2\pi i\eta x}\mathcal{Z}f(x+\xi,\omega+\eta).$
If we consider the Zak transform of a Fourier transform of a function, then we obtain
(2.28) $\mathcal{Z}\widehat{f}(x,\omega)=e^{2\pi ix\omega}\mathcal{Z}f(\omega,-x).$
This follows directly from the Poisson summation formula, which, in its standard form, is
(2.29) $\sum_{k\in{\mathbb{Z}}}f(k)=\sum_{l\in{\mathbb{Z}}}\widehat{f}(l).$
As the Fourier transform intertwines translation and modulation, precisely
(2.30) $\mathcal{F}T_{x}=M_{-x}\mathcal{F}\quad\text{ and }\quad\mathcal{F}M_{\omega}=T_{\omega}\mathcal{F},$
the phase factor in (2.27) appears then due to the commutation relation (2.9). We also note that the Zak transform extends to a unitary operator from $L^{2}({\mathbb{R}})$ onto $L^{2}([0,1]^{2})$. For the details of this fact we refer to [5, Thm. 8.2.3].
We close the survey on the Zak transform with listing zeros which occur due to the parity of $f$. For more details we refer to [9, § 5] or [8, § 4]. We have (see Fig. 1), for suitable $f$,
  • if $f(t)=f(-t)$, so $f$ is an even function, then
    (2.31) $\mathcal{Z}f\left(\tfrac{1}{2}+k,\tfrac{1}{2}+l\right)=0,\quad(k,l)\in{\mathbb{Z}}^{2},$
  • if $f(t)=-f(-t)$, so $f$ is an odd function, then
    (2.32) $\mathcal{Z}f\left(k,l\right)=\mathcal{Z}f\left(\tfrac{1}{2}+k,l\right)=\mathcal{Z}f\left(k,\tfrac{1}{2}+l\right)=0,\quad(k,l)\in{\mathbb{Z}}^{2}.$
The condition that $f$ is suitable cannot be dropped here. It implies that $\mathcal{Z}f$ is continuous. A remarkable property of the Zak transform is that, if $\mathcal{Z}f$ is continuous, then it has at least a zero in $[0,1)^{2}$. The indicator function $\chi_{[0,1]}$ is not suitable. It is not continuous, hence, in neither of the mentioned suitable function spaces and $|\mathcal{Z}\chi_{[0,1]}|=1$ almost everywhere.
Figure 1. Zeros of $\mathcal{Z}f(x,\omega)$ of a suitable even (left) and odd (right) function in $[0,1)^{2}$, obtained by parity. Further zeros may exist, depending on $f$.
3. Results
Theorem 3.1.
Let $\{z_{1},\ldots,z_{M}\}\subset[0,1)^{2}$ be distinct zeros of the Zak transform of $g$, i.e., $\mathcal{Z}g(z_{m})=0$, $m=1,\ldots,M$. Then, the multi-window Gabor system $\mathfrak{G}_{M}(\pi(z_{m})g,{\mathbb{Z}}^{2})$ is not a frame. Equivalently, the Gabor system $\mathcal{G}\left(g,\bigcup_{m=1}^{M}({\mathbb{Z}}^{2}+z_{m})\right)$ is not a frame.
Proof.
To keep this article self-contained, we repeat a standard proof, as given in the book of Gröchenig [5, Chap. 8.3], that the Zak transform diagonalizes the frame operator. More precisely, for a mulit-window Gabor system we will show (cf. [5, eq. (8.18)]) that
(3.1) $\mathcal{Z}\sum_{m=1}^{M}S_{\mathcal{G}(g_{m},{\mathbb{Z}}^{2})}f=\left(\sum_{m=1}^{M}|\mathcal{Z}g_{m}|^{2}\right)\mathcal{Z}f,\quad f\in L^{2}({\mathbb{R}}).$
To keep notation simple, we carry out the calculations for the case of a single window, as
(3.2) $\mathcal{Z}\left(\sum_{m=1}^{M}S_{\mathcal{G}(g_{m},{\mathbb{Z}}^{2})}f\right)=\sum_{m=1}^{M}\left(\mathcal{Z}S_{\mathcal{G}(g_{m},{\mathbb{Z}}^{2})}f\right),$
We start with combining (2.27) and (2.26) to conclude that
(3.3) $\mathcal{Z}(M_{l}T_{k}g_{m})(x,\omega)=e^{-2\pi ilx}\mathcal{Z}g_{m}(x+k,\omega+l)=e^{2\pi i(k\omega-lx)}\mathcal{Z}g_{m}(x,\omega),\quad k,l\in{\mathbb{Z}}.$
Next, using that $\mathcal{Z}:L^{2}({\mathbb{R}})\to L^{2}([0,1]^{2})$ is unitary, we compute
(3.4) $\mathcal{Z}(S_{\mathcal{G}(g_{m},{\mathbb{Z}}^{2})}f)(x,\omega)$ $=\sum_{(k,l)\in{\mathbb{Z}}^{2}}\langle f,M_{l}T_{k}g_{m}\rangle_{L^{2}({\mathbb{R}})}\,\mathcal{Z}(M_{l}T_{k}g_{m})(x,\omega)$
(3.5) $=\left(\sum_{(k,l)\in{\mathbb{Z}}^{2}}\langle\mathcal{Z}f,\mathcal{Z}(M_{l}T_{k}g_{m})\rangle_{L^{2}([0,1]^{2})}\,e^{2\pi i(k\omega-lx)}\right)\,\mathcal{Z}g_{m}(x,\omega)$
We have a look at the inner product. Using (3.3) we obtain
(3.6) $\langle\mathcal{Z}f,\mathcal{Z}(M_{l}T_{k}g_{m})\rangle_{L^{2}([0,1]^{2})}=\iint_{[0,1]^{2}}\mathcal{Z}f(x,\omega)\overline{\mathcal{Z}g_{m}(x,\omega)}e^{-2\pi i(k\omega-lx)}\,d(x,\omega).$
These are Fourier coefficients of $\mathcal{Z}f\overline{\mathcal{Z}g_{m}}$ and the whole expression in the parentheses in (3.5) is simply a Fourier series of this expression. Thus,
(3.7) $\sum_{m=1}^{M}\mathcal{Z}S_{\mathcal{G}(g_{m},{\mathbb{Z}}^{2})}f\left(\sum_{m=1}^{M}|\mathcal{Z}g_{m}|^{2}\right)\mathcal{Z}f\quad\Longleftrightarrow\quad\sum_{m=1}^{M}\mathcal{Z}S_{\mathcal{G}(g_{m},{\mathbb{Z}}^{2})}\mathcal{Z}^{-1}=\sum_{m=1}^{M}|\mathcal{Z}g_{m}|^{2}$
This was the first part of the proof. We also see that $\sum_{m=1}^{M}\mathcal{Z}S_{\mathcal{G}(g,{\mathbb{Z}}^{2})}\mathcal{Z}^{-1}$ is a multiplication operator which has the same spectrum as the frame operator. A multiplication operator $f\mapsto Mf$ is bounded if and only if $M\in L^{\infty}$ and invertible if and only if $M^{-1}\in L^{\infty}$. Hence,
(3.8) $\bigcup_{m=1}^{M}\mathcal{G}\left(\pi(z_{m})g,{\mathbb{Z}}^{2}\right)\text{ isa frame}\quad\Longleftrightarrow\quad 0<A\leq\sum_{m=1}^{M}|\mathcal{Z}(\pi(z_{m})g)(x,\omega)|^{2}\leq B<\infty.$
By using (2.27) again, we see that
(3.9) $\sum_{m=1}^{M}|\mathcal{Z}(\pi(z_{m})g)(z)|^{2}=\sum_{m=1}^{M}|\mathcal{Z}(g)(z+z_{m})|^{2}\quad z=(x,\omega).$
By assumption $\mathcal{Z}g(z_{m})=0$ for $m=1,\ldots,M$. Thus, by choosing $z=0$, we see that
(3.10) $\min_{z\in[0,1)^{2}}\sum_{m=1}^{M}|\mathcal{Z}(g)(z+z_{m})|^{2}=\sum_{m=1}^{M}|\mathcal{Z}(g)(z_{m})|^{2}=0.$
forcing the lower frame bound to vanish and the result is proved. ∎
At this point we remark, that there are several versions of the Zak transform. For example, in [8] or [10] the following version is used (re-call $\mathcal{D}_{a}$ from (2.3))
(3.11) $\widetilde{\mathcal{Z}}_{a}f(x,\omega)=\sqrt{a}\sum_{k\in{\mathbb{Z}}}f(a(k-x))e^{2\pi ik\omega}=\sum_{k\in{\mathbb{Z}}}M_{\omega}T_{x}\mathcal{D}_{a}^{-1}f(k)=\mathcal{Z}(\mathcal{D}_{a}^{-1}f)(x,\omega).$
This can be used to study Gabor systems of the form $\mathcal{G}(\mathcal{D}_{a}^{-1}g,{\mathbb{Z}}^{2})$. By unitary equivalence, this is the same as studying properties of $\mathcal{G}(g,D_{a}{\mathbb{Z}}^{2})$, so a Gabor system with window g over the rectangular lattice $a{\mathbb{Z}}\times\frac{1}{a}{\mathbb{Z}}$. On the other hand, there is the following version, which is for example presented in [9]. Setting $z=(x,\omega)\in{\mathbb{R}}^{2}$, we can write this version as
(3.12) $\mathcal{Z}_{a}f(z)$ $=\sqrt{a}\sum_{k\in{\mathbb{Z}}}f(ak-x)e^{2\pi iak\omega}$
(3.13) $=\sum_{k\in{\mathbb{Z}}}\mathcal{D}_{a}^{-1}\pi(z)f(k)=\sum_{k\in{\mathbb{Z}}}\pi(D_{a}^{-1}z)\mathcal{D}_{a}^{-1}f(k)$
(3.14) $=\mathcal{Z}(\mathcal{D}_{a}^{-1}f)(D_{a}^{-1}z).$
The essential properties from $\mathcal{Z}$ carry over to $\mathcal{Z}_{a}$, but the main difference is that $\mathcal{Z}_{a}$ is quasi-periodic with respect to the lattice $a{\mathbb{Z}}\times\frac{1}{a}{\mathbb{Z}}$. Hence, we have that $\mathcal{Z}_{a}$ is determined by its values on $[0,a)\times[0,\frac{1}{a})$ and it is a unitary operator from $L^{2}({\mathbb{R}})$ onto $L^{2}(D_{a}[0,1]^{2})$. This version of the Zak transform, combined with the techniques presented in the proof above, has long been used to study Gabor systems over rectangular lattices of the form $a{\mathbb{Z}}\times b{\mathbb{Z}}$ with $(ab)^{-1}\in{\mathbb{N}}$. The technique has also been expanded to the case that $(ab)^{-1}=\frac{p}{q}\in{\mathbb{Q}}$, leading to determining whether certain $q\times q$ matrices have full rank (see [5, Chap. 8.3]). Having a rectangular lattice as an index set is quite restrictive, as the shifts $\pi(\xi_{n})$ must possess very restricted symmetries. This is discussed to some extent in [8, § 3]. The freedom of allowing arbitrary periodic configurations leads to the following result.
Corollary 3.2.
Let $g\in M^{1}({\mathbb{R}})$ and assume $\mathcal{Z}g$ only has $N$ separated zeros in $[0,1)^{2}$. Pick $K>1$ points $\{\xi_{1},\ldots,\xi_{K}\}$ from $[0,1)^{2}$ i.i.d. randomly, then the Gabor system
(3.15) $\mathcal{G}\left(g,\bigcup_{k=1}^{K}({\mathbb{Z}}^{2}+\xi_{k})\right)\text{ isa frame with probability 1.}$.
If $K>N$, then $\mathcal{G}(g,\bigcup_{k=1}^{K}({\mathbb{Z}}^{2}+\xi_{k}))$ is a frame for any choice of $K$ distinct points in $[0,1)^{2}$.
Proof.
The fact that $g\in M^{1}({\mathbb{R}})$ guarantees that we have a finite upper frame bound. As the zero set of $\mathcal{Z}g$ is discrete and finite by assumption, the probability that $K$ i.i.d. randomly picked picking points from $[0,1)^{2}$ all belong to the zero set of $\mathcal{Z}g$ is zero. ∎
Corollary 3.3.
For the $n$-th Hermite function $h_{n}$, the Zak transform $\mathcal{Z}h_{n}$ can have at most $n+1$ zeros $\{z_{1},\ldots,z_{n+1}\}\subset[0,1)^{2}$ such that
(3.16) $\Gamma_{n+1}({\mathbb{Z}}^{2},\{z_{\ell}\}_{\ell=1}^{n+1})=\bigcup_{\ell=1}^{n+1}({\mathbb{Z}}^{2}+z_{\ell})\text{ is a lattice.}$.
Proof.
Due to the results of Gröchenig and Lyubarskii [6][7] we know that if we have a Gabor system over a lattice $\Lambda$ with ${\rm vol}({\mathbb{R}}^{2}/\Lambda)<1/(n+1)$, then $\mathcal{G}(h_{n},\Lambda)$ is a frame. ∎
Note, that this does not exclude the option that $\mathcal{Z}h_{n}$ has more than $n+1$ zeros in $[0,1)^{2}$.
Corollary 3.4.
There exist Gabor systems with Hermite functions of order $n\geq 1$ over a periodic index set of density strictly greater than $n+1$ which are not frames.
The proof of Corollary 3.4 is by example. We start with the first Hermite function.
Example 1. Consider the Zak transform of $h_{1}$. Since $h_{1}$ is odd, we know that
(3.17) $\mathcal{Z}h_{1}(0,0)=\mathcal{Z}h_{1}\left(\tfrac{1}{2},0\right)=\mathcal{Z}h_{1}\left(0,\tfrac{1}{2}\right)=0.$
In particular, the Zak transform $\mathcal{Z}h_{1}$ has (at least) 3 zeros in $[0,1)^{2}$. Thus, there exists a periodic configuration $\Gamma$ of density 3 such that $\mathcal{G}(h_{1},\Gamma)$ is not a frame. As we know the exact location of the 3 zeros, we know $\Gamma$, which we depict in Figure 2. Concretely, the Gabor system $\mathcal{G}\left(h_{1},{\mathbb{Z}}^{2}\cup(({\mathbb{Z}}+\frac{1}{2})\times{\mathbb{Z}})\cup({\mathbb{Z}}\times({\mathbb{Z}}+\frac{1}{2}))\right)$ is not a frame.
Figure 2. The set $\Gamma=\bigcup_{m=1}^{3}({\mathbb{Z}}^{2}+z_{m})$, $z_{m}\in\{(0,0),(\frac{1}{2},0),(0,\frac{1}{2})\}$ and its decomposition into 3 relatively shifted copies of the integer lattice ${\mathbb{Z}}^{2}$.
Example 2 We consider the third Hermite function $h_{3}$. In [10], Lemvig lists zeros of $\mathcal{Z}\mathcal{D}_{\sqrt{a}}^{-1}h_{4\ell+3}$, $\ell\in{\mathbb{N}}_{0}$, $a\in\{\frac{1}{4},\frac{1}{3},\frac{1}{2},2,3,4\}$. These were earlier discovered by Boon, Zak, and Zucker [1] in a different context for $\mathcal{Z}_{a}h_{4\ell+3}$. Due to a different normalization of the Fourier transform $\mathcal{Z}_{a}$ is defined slightly different in [1] compared to (3.12). We have the following known zeros (further zeros may exist), as listed in [1] (subject to our normalization):
$\bullet$ $\mathcal{Z}_{\sqrt{2}}h_{4\ell+3}(z)=0$, $z\in\left\{(0,0),\left(\frac{\sqrt{2}}{2},0\right),\left(0,\frac{1}{2\sqrt{2}}\right),\left(\frac{\sqrt{2}}{4},\frac{1}{2\sqrt{2}}\right),\left(\frac{3\sqrt{2}}{4},\frac{1}{2\sqrt{2}}\right)\right\}+D_{\sqrt{2}}{\mathbb{Z}}^{2}$.
The location of the zeros in $[0,\sqrt{2})\times[0,\frac{1}{\sqrt{2}})$ is depicted in Figure 3.
Figure 3. Known zeros of $\mathcal{Z}_{\sqrt{2}}h_{4\ell+3}$ in the fundamental cell of $\sqrt{2}{\mathbb{Z}}\times\frac{1}{\sqrt{2}}{\mathbb{Z}}$.
$\bullet$ $\mathcal{Z}_{\sqrt{3}}h_{4\ell+3}(z)=0$, $z\in\left\{(0,0),\left(0,\frac{1}{2\sqrt{3}}\right),\left(\frac{\sqrt{3}}{2},0\right),\left(\frac{\sqrt{3}}{3},0\right),\left(\frac{2\sqrt{3}}{3},0\right)\right\}+D_{\sqrt{3}}{\mathbb{Z}}^{2}$.
The location of the zeros in $[0,\sqrt{3})\times[0,\frac{1}{\sqrt{3}})$ is depicted in Figure 4.
Figure 4. Known zeros for $\mathcal{Z}_{\sqrt{3}}h_{4\ell+3}$ in the fundamental cell of $\sqrt{3}{\mathbb{Z}}\times\frac{1}{\sqrt{3}}{\mathbb{Z}}$.
$\bullet$ $\mathcal{Z}_{2}h_{4\ell+3}(z)=0$, $z\in\left\{(0,0),\left(0,\frac{1}{4}\right),\left(1,0\right),\left(\frac{1}{2},0\right),\left(\frac{3}{2},0\right)\right\}+D_{2}{\mathbb{Z}}^{2}$.
The location of the zeros in $[0,\sqrt{3})\times[0,\frac{1}{\sqrt{3}})$ is depicted in Figure 4.
Figure 5. Known zeros for $\mathcal{Z}_{2}h_{4\ell+3}$ in the fundamental cell of $2{\mathbb{Z}}\times\frac{1}{2}{\mathbb{Z}}$.
Note that the first three zeros in any of the above lists are due to the parity of $h_{4\ell+3}$. The cases for $\frac{1}{\sqrt{a}}$ are covered by an application of the Poisson summation formula, similarly to (2.28), which results in simply rotating the rectangle in Figure 3 by 90 degrees.
Thus, specifically picking $h_{3}$ (i.e., $\ell=0$) as window function, we are able to design a wealth of Gabor systems over periodic configurations of density 5 which are not frames.